A guy is 62 1/2 inches tall his daughter is 57 1/9 inches tall how much taller is the guy
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

62-57=5
1/2-1/9 convert to 18ths
9/18-2/18=7/18
5 7/18 inches taller depending on hairstyle and heals.

Answer by stanbon(75887)

A father is 62 1/2 inches tall; his daughter is 57 1/9 inches tall.
how much taller is the father?
---------------
62 1/2 - 57 1/9
----
= (62-57)+(1/2 - 1/9)
----
= 5 + (9-2)/18
----
= 5 7/18 inches
================
Cheers,
Stan H.
================